OpenStudy (anonymous):

abcd is a rectangle find the length of each diagonal ac=2(5a+1) bd=2(a+1)

OpenStudy (anonymous):

the diagonals in a rectangle are equal=>2(5a+1)=2(a+1) =>10a+2=2a+2 =>a=0 =>AC=BD=2 units of length